# EntroPE: Entropy-Guided Dynamic Patch Encoder for Time Series Forecasting

### v1

@@ A.1 Baseline Details
TimeKAN (Huang et al., 2025b) Kernel Attn. Network Kernelized attention distinct from Transformer/MLP.
@@
CALF (Liu et al., 2025) LLM-aligned / Cross-modal Cross-modal match, feature regularization, output consistency.
@@ A.2 Dataset Details
Exchange Rate 8 (5120, 665, 1422) Daily Foreign Exchange Market
@@
  • Exchange Rate: This dataset contains daily exchange rates of the US dollar against eight different currencies, spanning from 1990 to 2016. It represents financial time series forecasting challenges with inherent volatility and complex market dynamics.
@@
A.3 Implementation Details
@@
Embedding dimension (n_embdn\_embd) 8
@@
Hyperparameter ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity Exchange Rate
dim 8 8 16 16 16 32 8
heads 2 2 2 4 2 4 2
layers 1 2 1 1 2 2 1
max patch length 24 24 24 24 24 24 16
batch_size 64 64 32 32 128 32 32
learning_rate 0.001 0.01 0.01 0.001 0.01 0.01 0.01
dropout 0.05 0.1 0.1 0.1 0.2 0.1 0.2
monotonicity 1 1 1 1 1 1 1
threshold ($\theta$) 3 3.5 3.5 3 3 3.5 3.5
threshold_add ($\gamma$) 0.25 0.2 0.3 0.25 0.35 0.3 0.3
train_epochs 20 20 20 20 20 20 20
patience 7 7 7 7 7 7 7
A.4 Extended Results
A.4.1 Full Results (Long-term forecasting). Input length 96.
@@ A.4 Extended Results
We adopted the training criteria established by TimeKAN for fair comparison across all baseline methods. The results for AutoFormer, FEDformer, TimesNet, PatchTST, Time-FFM, iTransformer, and TimeMixer were directly extracted from the TimeKAN paper to ensure consistency in experimental conditions. For HDMixer, TimeBase, CALF, and FilterTS, we executed the original implementations following identical training criteria. Specifically for FilterTS, we addressed the challenge of multiple configuration options by running experiments across all different settings proposed by the authors for each forecasting horizon. We then selected the unified setting that achieved the lowest Mean Squared Error (MSE) and Mean Absolute Error (MAE) performance (detailed results are provided in Section A.4.2). LangTime baseline results were obtained directly by running the author’s implementation to maintain consistency with the their reported performance. This methodology ensures that all comparative results are obtained under equivalent experimental conditions, providing a fair and rigorous evaluation framework for our proposed approach. See Tab 7 and Tab. 8 for full results.
@@
Dataset T EntroPE TimeKAN HDMixer TimeBase CALF FilterTS LangTime TimeMixer Time-FFM
@@
ETTm1 96 0.317 0.354 0.327 0.365 0.340 0.369 0.373 0.388 0.319 0.348 0.323 0.362 0.319 0.348 0.317 0.356 0.336 0.369
192 0.360 0.379 0.354 0.380 0.379 0.385 0.411 0.409 0.375 0.376 0.364 0.382 0.368 0.375 0.367 0.384 0.378 0.389
336 0.388 0.399 0.386 0.403 0.398 0.409 0.436 0.421 0.410 0.399 0.396 0.404 0.413 0.402 0.391 0.406 0.411 0.410
720 0.446 0.433 0.452 0.442 0.467 0.443 0.503 0.461 0.478 0.439 0.462 0.441 0.487 0.439 0.454 0.441 0.469 0.441
Avg. 0.378 0.391 0.380 0.398 0.396 0.402 0.431 0.420 0.396 0.391 0.386 0.397 0.397 0.392 0.382 0.397 0.399 0.402
ETTm2 96 0.180 0.265 0.175 0.258 0.183 0.266 0.188 0.271 0.174 0.252 0.174 0.256 0.188 0.258 0.175 0.257 0.181 0.267
192 0.241 0.311 0.241 0.301 0.246 0.306 0.251 0.309 0.241 0.297 0.240 0.299 0.245 0.297 0.240 0.302 0.247 0.308
336 0.313 0.352 0.308 0.349 0.307 0.347 0.311 0.346 0.304 0.338 0.301 0.339 0.301 0.336 0.303 0.343 0.309 0.347
720 0.411 0.413 0.417 0.414 0.408 0.404 0.411 0.401 0.402 0.395 0.399 0.399 0.402 0.393 0.392 0.396 0.406 0.404
Avg. 0.286 0.335 0.285 0.331 0.286 0.331 0.290 0.332 0.280 0.321 0.279 0.323 0.284 0.321 0.279 0.324 0.286 0.332
ETTh1 96 0.375 0.399 0.367 0.395 0.387 0.401 0.399 0.392 0.377 0.395 0.377 0.391 0.391 0.388 0.385 0.402 0.385 0.400
192 0.423 0.425 0.416 0.424 0.441 0.428 0.455 0.423 0.429 0.427 0.431 0.423 0.429 0.419 0.443 0.430 0.439 0.430
336 0.429 0.432 0.446 0.436 0.452 0.433 0.501 0.443 0.475 0.449 0.479 0.448 0.462 0.440 0.512 0.470 0.480 0.449
720 0.439 0.454 0.441 0.453 0.513 0.485 0.498 0.458 0.482 0.470 0.471 0.466 0.458 0.445 0.497 0.476 0.462 0.456
@@
ETTh2 96 0.281 0.336 0.290 0.339 0.289 0.336 0.338 0.376 0.288 0.336 0.293 0.344 0.299 0.336 0.289 0.342 0.301 0.351
192 0.371 0.393 0.386 0.399 0.386 0.397 0.402 0.405 0.368 0.386 0.374 0.389 0.374 0.382 0.378 0.397 0.378 0.397
336 0.392 0.394 0.441 0.447 0.438 0.442 0.437 0.440 0.415 0.423 0.411 0.423 0.410 0.418 0.432 0.434 0.422 0.431
720 0.421 0.427 0.450 0.458 0.421 0.454 0.460 0.477 0.417 0.435 0.423 0.441 0.418 0.426 0.464 0.464 0.427 0.444
Avg. 0.366 0.387 0.391 0.410 0.384 0.407 0.409 0.425 0.372 0.395 0.375 0.399 0.375 0.392 0.390 0.409 0.382 0.406
Weather 96 0.164 0.211 0.163 0.209 0.175 0.223 0.170 0.215 0.164 0.205 0.161 0.208 0.178 0.204 0.163 0.209 0.191 0.230
192 0.210 0.252 0.208 0.251 0.226 0.265 0.216 0.256 0.213 0.252 0.226 0.264 0.211 0.249 0.211 0.254 0.236 0.267
336 0.256 0.290 0.263 0.290 0.264 0.301 0.272 0.297 0.270 0.292 0.279 0.304 0.269 0.292 0.263 0.293 0.289 0.303
720 0.339 0.342 0.341 0.341 0.348 0.349 0.351 0.348 0.352 0.346 0.345 0.344 0.351 0.347 0.344 0.348 0.362 0.350
Avg. 0.242 0.273 0.244 0.273 0.253 0.285 0.252 0.279 0.250 0.274 0.253 0.280 0.252 0.273 0.245 0.276 0.270 0.288
Electricity 96 0.163 0.252 0.174 0.266 0.180 0.271 0.212 0.279 0.145 0.239 0.153 0.247 0.181 0.266 0.153 0.245 0.198 0.282
192 0.177 0.268 0.182 0.273 0.184 0.275 0.209 0.281 0.162 0.253 0.168 0.260 0.185 0.273 0.166 0.257 0.199 0.285
336 0.194 0.284 0.197 0.286 0.207 0.299 0.222 0.295 0.177 0.268 0.187 0.278 0.198 0.281 0.185 0.275 0.212 0.298
720 0.235 0.321 0.236 0.320 0.249 0.335 0.264 0.327 0.222 0.304 0.227 0.313 0.241 0.320 0.224 0.312 0.253 0.330
Avg. 0.182 0.271 0.197 0.286 0.205 0.295 0.227 0.296 0.177 0.266 0.184 0.275 0.201 0.285 0.182 0.272 0.216 0.299
Exchange 96 0.083 0.140 0.084 0.201 0.089 0.201 0.087 0.205 0.081 0.201
192 0.177 0.235 0.173 0.295 0.175 0.298 0.179 0.300 0.168 0.293
336 0.299 0.371 0.316 0.407 0.329 0.409 0.333 0.417 0.299 0.396
720 0.761 0.869 0.827 0.685 0.852 0.690 0.912 0.719 0.805 0.674
Avg. 0.331 0.386 0.355 0.400 0.361 0.400 0.378 0.410 0.338 0.391

### v2

@@ D.1 Baseline Details
TimeMosaic (Ding et al., 2025) Transformer patch granularity based on temporal heterogeneity and motif reuse.
@@
TSLANet (Eldele et al., 2024) CNN + Lightweight Lightweight and pretraining module for enhanced feature representation
@@
CALF (Liu et al., 2025b) LLM-aligned / Cross-modal Cross-modal match, feature regularization, output consistency.
@@ D.2 Dataset Details
D.3 Classification Dataset
For classification we used 10 UEA datasets, following the evaluation and preprocessing protocol of TSLANet (Eldele et al., 2024).
UEA datasets: We incorporate datasets from the University of East Anglia (UEA) Time Series Classification repository, which is renowned for its rich collection of multivariate time series datasets. We were able to preprocess 10 datasets, each offering a multidimensional perspective on time series analysis across various real-world scenarios, such as human activity recognition, sensor data interpretation, and complex system monitoring. More details about the UEA datasets can be found in https://www.timeseriesclassification.com/.
D.4 Anomaly Detection Dataset
For Anomaly Detection also we followed the evaluation and preprocessing protocol of TSLANet. In our study, we assess the performance of our model using five benchmark datasets, each representing a distinct application area, to demonstrate its effectiveness in detecting anomalies in diverse settings:
  • SMD (Server Machine Dataset): Utilized for server monitoring, the SMD dataset comprises multivariate time series data collected from servers and aims to identify unusual server behaviors that could indicate failures or security issues.
  • MSL (Mars Science Laboratory): This dataset contains telemetry data from the Mars Science Laboratory rover, focusing on space exploration applications. Anomaly detection in this context is crucial for identifying potential issues with spacecraft systems based on their operational data.
  • SMAP (Soil Moisture Active Passive): Related to earth observations, the SMAP dataset includes soil moisture measurements intended for environmental monitoring. Detecting anomalies in soil moisture can provide insights into environmental conditions and potential agricultural impacts.
  • SWaT (Secure Water Treatment): In the domain of water treatment security, the SWaT dataset consists of data from a water treatment testbed, simulating the operational data of water treatment plants. Anomaly detection here is vital for ensuring the safety and security of water treatment processes.
  • PSM (Pump Sensor Monitoring): Focused on industrial pump sensors, the PSM dataset gathers sensor data from pumps in industrial settings. Anomalies in this dataset can indicate equipment malfunctions or the need for maintenance, critical for preventing industrial accidents.
The detailed characteristics of these datasets is presented in Table 7.
Table 7: Summary of datasets used for anomaly detection evaluation. Dataset sizes represent the number of temporal observations in each partition (Training, Validation, Test).
Dataset Dim Dataset Size (Train, Val, Test) Length Domain Information
SMD 38 (566724, 141681, 708420) 100 Server Machine
MSL 55 (44653, 11664, 73729) 100 Spacecraft
SMAP 25 (108146, 27037, 427617) 100 Spacecraft
SWaT 51 (396000, 99000, 449919) 100 Infrastructure
PSM 25 (105984, 26497, 87841) 100 Server Machine
D.5 Implementation Details
@@
Embedding dimension (n_embdn\_embd) 16
@@
Hyperparameter ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity
dim 8 8 16 16 16 32
heads 2 2 2 4 2 4
layers 1 2 1 1 2 2
max patch length 24 24 24 24 24 24
batch_size 64 64 32 32 128 32
learning_rate 0.001 0.01 0.01 0.001 0.01 0.01
dropout 0.05 0.1 0.1 0.1 0.2 0.1
threshold % ($\theta$) 3 3.5 3.5 3 3 3.5
train_epochs 20 20 20 20 20 20
Appendix E Extended Experimental Results
E.1 Time Series Classification
To demonstrate the generalizability of EntroPE beyond forecasting, we evaluate on 10 UEA time series classification datasets. We replace the forecasting head with a classification head while keeping the core architecture unchanged. Table 10 presents results compared to recent baselines.
Table 10: Classification accuracy (%) on UEA benchmark datasets. EntroPE achieves competitive or superior performance across diverse classification tasks.
Dataset EntroPE TSLANet GPT4TS TimesNet ROCKET CrossF. PatchTST MLP TS-TCC
EthanolConcentration 29.1 30.42 25.48 27.73 42.58 34.98 28.90 33.46 32.32
FaceDetection 69.1 66.77 65.58 67.47 64.70 66.17 68.96 67.42 63.05
Handwriting 57.4 57.88 3.76 26.18 48.47 26.24 26.00 22.47 47.76
Heartbeat 77.9 77.56 36.59 74.48 69.76 76.59 76.59 73.17 77.07
JapaneseVowels 99.0 99.19 98.11 97.83 95.68 98.92 98.65 97.84 97.30
PEMS-SF 88.6 83.82 87.28 88.13 75.10 82.08 88.44 82.08 86.71
SelfRegulationSCP1 92.9 91.81 91.47 77.43 84.64 92.49 89.76 88.40 91.13
SelfRegulationSCP2 63.1 61.67 51.67 52.84 54.44 53.33 54.44 51.67 53.89
SpokenArabicDigits 99.8 99.91 99.36 98.36 99.20 96.41 99.68 96.68 99.77
UWaveGestureLibrary 94.4 91.25 84.38 83.13 94.40 81.56 80.00 81.88 86.25
Average 77.13 76.03 64.37 69.36 72.90 70.88 71.14 69.51 73.53
EntroPE achieves an average accuracy of 77.13%, outperforming TSLANet (76.03%), TimesNet (69.36%), PatchTST (71.14%), and other competitive baselines. This demonstrates that entropy-guided dynamic patching effectively captures discriminative temporal patterns for classification tasks.
E.2 Anomaly Detection
We further evaluate EntroPE on five standard anomaly detection benchmarks (SMD, MSL, SMAP, SWaT, PSM) by replacing the forecasting head with an anomaly scoring mechanism. Table 11 shows F1 scores compared to recent methods.
Table 11: Anomaly detection F1 scores (%) on benchmark datasets. EntroPE demonstrates strong performance across diverse anomaly detection scenarios.
Dataset EntroPE TSLANet GPT4TS TimesNet PatchTST ETSformer FEDformer LightTS DLinear
SMD 87.1 87.91 86.89 84.61 84.62 83.13 85.08 82.53 77.10
MSL 85.3 83.32 82.45 81.84 78.70 85.03 78.57 78.95 84.88
SMAP 74.2 75.96 72.88 69.39 68.82 69.50 70.76 69.21 69.26
SWaT 94.3 92.80 94.23 93.02 85.72 84.91 93.19 93.33 87.52
PSM 97.7 97.73 97.13 97.34 96.08 91.76 97.23 97.15 93.55
Average 87.72 87.54 86.72 85.24 82.79 82.87 84.97 84.23 82.46
EntroPE attains an average F1 score of 87.72%, competitive with or better than TSLANet (87.54%), GPT4TS (86.72%), TimesNet (85.24%), and other baselines. These results confirm that the entropy-guided patching mechanism generalizes beyond forecasting to anomaly detection tasks.
E.3 Comparison with Large-Scale Foundation Models
We compare EntroPE with recent large-scale time series foundation models including MOIRAI (Small, Base, Large), Sundial (Small, Base, Large), and TimeMoE (Base, Large, XL). Table 12 presents results on ETT and related datasets.
Table 12: Comparison with large-scale foundation models (MSE / MAE). EntroPE achieves competitive performance with significantly fewer parameters.
Dataset EntroPE MOIRAI_S MOIRAI_B MOIRAI_L Sundial_S Sundial_B Sundial_L TimeMoE_B TimeMoE_L TimeMoE_XL
ETTm1 0.378/0.391 0.400/0.424 0.434/0.439 0.510/0.468 0.354/0.388 0.336/0.377 0.331/0.369 0.394/0.415 0.376/0.405 0.341/0.382
ETTm2 0.286/0.335 0.311/0.348 0.279/0.330 0.282/0.330 0.265/0.324 0.258/0.320 0.254/0.315 0.317/0.365 0.316/0.361 0.288/0.344
ETTh1 0.416/0.425 0.400/0.423 0.434/0.439 0.510/0.468 0.390/0.418 0.411/0.434 0.395/0.420 0.400/0.424 0.394/0.409 0.412/0.426
ETTh2 0.366/0.387 0.348/0.374 0.354/0.376 0.359/0.371 0.340/0.387 0.333/0.387 0.334/0.387 0.366/0.404 0.405/0.415 0.371/0.399
Electricity 0.182/0.271 0.238/0.303 0.219/0.283 0.229/0.281 0.169/0.265 0.169/0.265 0.166/0.262
Weather 0.242/0.273 0.233/0.271 0.234/0.270 0.238/0.275 0.233/0.271 0.234/0.270 0.238/0.275 0.265/0.297 0.270/0.300 0.256/0.288
Despite having orders of magnitude fewer parameters (EntroPE: $\sim$0.1-0.2M vs. foundation models: 10M-500M+), EntroPE achieves competitive or superior performance on several datasets, demonstrating the effectiveness of entropy-guided dynamic patching as a lightweight alternative to massive pre-trained models.
E.4 Comprehensive Efficiency Comparison
Table 13 presents a detailed comparison of parameters and MACs (Multiply-Accumulate Operations) across multiple datasets for EntroPE and recent baselines.
Table 13: Comprehensive efficiency analysis: Parameters and MACs across datasets. EntroPE achieves competitive accuracy with significantly lower computational cost.
Model ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity
Params MACs Params MACs Params MACs Params MACs Params MACs Params MACs
TimeMixer 75.50K 20.37M 75.50K 20.37M 75.50K 20.37M 77.77K 24.18M 104.43K 82.62M 106.83K 1.26G
iTransformer 841.57K 77.46M 224.22K 19.86M 224.22K 19.86M 224.22K 19.86M 4.83M 1.16G 4.83M 16.29G
PatchTST 3.75M 5.90G 10.06M 17.66G 3.75M 5.90G 10.06M 17.66G 6.90M 35.30G 6.90M 539.38G
TimesNet 605.48K 18.13G 1.19M 36.28G 4.71M 144G 1.19M 36.28G 1.19M 36.28G 150.30M 4.61T
MICN 25.20M 71.95G 25.20M 71.95G 25.20M 71.95G 25.20M 71.95G 111.03M 295.07M 6.64M 19.5G
FiLM 12.58M 2.82G 12.58M 2.82G 12.58M 2.82G 12.58M 2.82G 12.58M 8.46G 12.58M 8.46G
FEDFormer 23.38M 24.96G 23.38M 24.96G 23.38M 24.96G 23.38M 24.96G 23.45M 25.23G 24.29M 30.89G
AutoFormer 10.54M 22.82G 10.54M 22.82G 10.54M 22.82G 10.54M 22.82G 10.61M 22.98G 12.14M 28.75G
EntroPE 95K 9M 418K 53M 192K 20M 95K 10M 195K 61M 199K 1G
[…]
E.5 EntroPE as Plug-in Replacement for PatchTST
To demonstrate the modularity and effectiveness of entropy-guided patching, we integrate EntroPE’s dynamic patcher and Fusion Decoder into PatchTST while keeping other components unchanged. Table 14 shows the performance improvement.
Table 14: EntroPE as plug-in replacement for PatchTST’s static patcher. Relative improvement (%) shows consistent gains across datasets.
Model ETTh1 ETTh2 ETTm1 ETTm2 Weather ECL
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
PatchTST 0.516 0.484 0.391 0.411 0.406 0.407 0.290 0.334 0.265 0.285 0.216 0.318
EntroPE 0.416 0.425 0.366 0.387 0.378 0.391 0.286 0.335 0.242 0.273 0.182 0.271
$\Delta$% Improvement 19% 12% 6% 6% 7% 4% 1% 0% 9% 4% 16% 15%
Replacing PatchTST’s static patcher with EntroPE’s entropy-guided dynamic patcher yields approximately 8-10% average MSE improvement across datasets, reduces parameters from $\sim$4M to 50K-1M range, and decreases training time. This demonstrates that EntroPE can serve as a drop-in replacement for static patching mechanisms in existing architectures.
E.6 Training Efficiency Analysis
We measure end-to-end training time per epoch to assess the computational overhead of dynamic patching. Table 15 shows that entropy-guided patching adds minimal overhead.
Table 15: Training efficiency comparison: Epoch duration (seconds) for static vs. dynamic patching.
Metric/Dataset ETTh1 ETTh2 ETTm1 ETTm2 Weather ECL
Epoch (dynamic) 3.2 3.5 17.2 20.8 54.5 181
Epoch (static) 3.1 3.4 17.0 20.5 54.1 180
Patching time 0.1 0.1 0.2 0.3 0.4 1
Overhead % 3% 3% 1% 1% 1% 1%
Dynamic patching increases epoch time by only 0.1-1.0 seconds (1-3% overhead), demonstrating that the entropy computation and boundary detection are computationally negligible compared to the transformer’s forward-backward passes.
E.6.1 Full Results (Long-term forecasting). Input length 96.
@@ E.6 Training Efficiency Analysis
We adopted the training criteria established by TimeKAN and TimesNet for fair comparison across all baseline methods. The results for AutoFormer, FEDformer, TimesNet, PatchTST, Time-FFM, iTransformer, and TimeMixer were directly extracted from the TimeKAN paper to ensure consistency in experimental conditions. For HDMixer, TimeBase, CALF, TSLANet, TimeMosaic and FilterTS we executed the original implementations following identical training criteria. Specifically for FilterTS, we addressed the challenge of multiple configuration options by running experiments across all different settings proposed by the authors for each forecasting horizon. We then selected the unified setting that achieved the lowest Mean Squared Error (MSE) and Mean Absolute Error (MAE) performance (detailed results are provided in Section E.6.2). LangTime baseline results were obtained directly by running the author’s implementation to maintain consistency with the their reported performance. This methodology ensures that all comparative results are obtained under equivalent experimental conditions, providing a fair and rigorous evaluation framework for our proposed approach. See Tab 16 and Tab. 17 for full results.
@@
Dataset T EntroPE TimeMosaic HDMixer TimeBase CALF FilterTS LangTime TimeMixer Time-FFM
@@
ETTm1 96 0.317 0.354 0.312 0.351 0.340 0.369 0.373 0.388 0.319 0.348 0.323 0.362 0.319 0.348 0.317 0.356 0.336 0.369
192 0.360 0.379 0.370 0.379 0.379 0.385 0.411 0.409 0.375 0.376 0.364 0.382 0.368 0.375 0.367 0.384 0.378 0.389
336 0.388 0.399 0.386 0.399 0.398 0.409 0.436 0.421 0.410 0.399 0.396 0.404 0.413 0.402 0.391 0.406 0.411 0.410
720 0.446 0.433 0.469 0.440 0.467 0.443 0.503 0.461 0.478 0.439 0.462 0.441 0.487 0.439 0.454 0.441 0.469 0.441
Avg. 0.378 0.391 0.388 0.392 0.396 0.402 0.431 0.420 0.396 0.391 0.386 0.397 0.397 0.392 0.382 0.397 0.399 0.402
ETTm2 96 0.180 0.265 0.177 0.254 0.183 0.266 0.188 0.271 0.174 0.252 0.174 0.256 0.188 0.258 0.175 0.257 0.181 0.267
192 0.241 0.311 0.240 0.295 0.246 0.306 0.251 0.309 0.241 0.297 0.240 0.299 0.245 0.297 0.240 0.302 0.247 0.308
336 0.313 0.352 0.298 0.336 0.307 0.347 0.311 0.346 0.304 0.338 0.301 0.339 0.301 0.336 0.303 0.343 0.309 0.347
720 0.411 0.413 0.397 0.391 0.408 0.404 0.411 0.401 0.402 0.395 0.399 0.399 0.402 0.393 0.392 0.396 0.406 0.404
Avg. 0.286 0.335 0.278 0.319 0.286 0.331 0.290 0.332 0.280 0.321 0.279 0.323 0.284 0.321 0.279 0.324 0.286 0.332
ETTh1 96 0.375 0.399 0.369 0.397 0.387 0.401 0.399 0.392 0.377 0.395 0.377 0.391 0.391 0.388 0.385 0.402 0.385 0.400
192 0.423 0.425 0.419 0.439 0.441 0.428 0.455 0.423 0.429 0.427 0.431 0.423 0.429 0.419 0.443 0.430 0.439 0.430
336 0.429 0.432 0.454 0.451 0.452 0.433 0.501 0.443 0.475 0.449 0.479 0.448 0.462 0.440 0.512 0.470 0.480 0.449
720 0.439 0.454 0.466 0.479 0.513 0.485 0.498 0.458 0.482 0.470 0.471 0.466 0.458 0.445 0.497 0.476 0.462 0.456
@@
ETTh2 96 0.281 0.336 0.292 0.341 0.289 0.336 0.338 0.376 0.288 0.336 0.293 0.344 0.299 0.336 0.289 0.342 0.301 0.351
192 0.371 0.393 0.367 0.391 0.386 0.397 0.402 0.405 0.368 0.386 0.374 0.389 0.374 0.382 0.378 0.397 0.378 0.397
336 0.392 0.394 0.417 0.427 0.438 0.442 0.437 0.440 0.415 0.423 0.411 0.423 0.410 0.418 0.432 0.434 0.422 0.431
720 0.421 0.427 0.398 0.436 0.421 0.454 0.460 0.477 0.417 0.435 0.423 0.441 0.418 0.426 0.464 0.464 0.427 0.444
Avg. 0.366 0.387 0.369 0.399 0.384 0.407 0.409 0.425 0.372 0.395 0.375 0.399 0.375 0.392 0.390 0.409 0.382 0.406
Weather 96 0.164 0.211 0.166 0.201 0.175 0.223 0.170 0.215 0.164 0.205 0.161 0.208 0.178 0.204 0.163 0.209 0.191 0.230
192 0.210 0.252 0.219 0.258 0.226 0.265 0.216 0.256 0.213 0.252 0.226 0.264 0.211 0.249 0.211 0.254 0.236 0.267
336 0.256 0.290 0.273 0.296 0.264 0.301 0.272 0.297 0.270 0.292 0.279 0.304 0.269 0.292 0.263 0.293 0.289 0.303
720 0.339 0.342 0.354 0.351 0.348 0.349 0.351 0.348 0.352 0.346 0.345 0.344 0.351 0.347 0.344 0.348 0.362 0.350
Avg. 0.242 0.273 0.253 0.277 0.253 0.285 0.252 0.279 0.250 0.274 0.253 0.280 0.252 0.273 0.245 0.276 0.270 0.288
Electricity 96 0.163 0.252 0.161 0.261 0.180 0.271 0.212 0.279 0.145 0.239 0.153 0.247 0.181 0.266 0.153 0.245 0.198 0.282
192 0.177 0.268 0.176 0.274 0.184 0.275 0.209 0.281 0.162 0.253 0.168 0.260 0.185 0.273 0.166 0.257 0.199 0.285
336 0.194 0.284 0.189 0.283 0.207 0.299 0.222 0.295 0.177 0.268 0.187 0.278 0.198 0.281 0.185 0.275 0.212 0.298
720 0.235 0.321 0.225 0.304 0.249 0.335 0.264 0.327 0.222 0.304 0.227 0.313 0.241 0.320 0.224 0.312 0.253 0.330
Avg. 0.182 0.271 0.188 0.281 0.205 0.295 0.227 0.296 0.177 0.266 0.184 0.275 0.201 0.285 0.182 0.272 0.216 0.299